# Hamiltonian and Liouvillian learning in weakly-dissipative quantum many-body systems

### v1

@@ Abstract
We discuss Hamiltonian and Liouvillian learning for analog quantum simulation from non-equilibrium quench dynamics in the limit of weakly dissipative many-body systems. We present various strategies to learn the operator content of the Hamiltonian and the Lindblad operators of the Liouvillian. We compare different ansätze based on an experimentally accessible ‘learning error’ which we consider as a function of the number of runs of the experiment. Initially, the learning error decreasing with the inverse square root of the number of runs, as the error in the reconstructed parameters is dominated by shot noise. Eventually the learning error remains constant, allowing us to recognize missing ansatz terms. A central aspect of our approach is to (re-)parametrize ansätze by introducing and varying the dependencies between parameters. This allows us to identify the relevant parameters of the system, thereby reducing the complexity of the learning task. Importantly, this (re-)parametrization relies solely on classical post-processing, which is compelling given the finite amount of data available from experiments. A distinguishing feature of our approach is the possibility to learn the Hamiltonian, without the necessity of learning the complete Liouvillian, thus further reducing the complexity of the learning task. We illustrate our method with two, experimentally relevant, spin models.
@@ I Introduction
An outstanding challenge in quantum simulation is the ability to predict properties of many-body observables with controlled error while scaling to a regime of potential quantum advantage [21, 22, 23, 24]. Given the increase of complexity of these systems, methods to characterize, and thus verify, the proper functioning of quantum simulators are required [25, 26, 27]. This includes verification, that the correct many-body Hamiltonians are being implemented, and a complete characterization of (weak) decoherence due to unwanted couplings to an environment or fluctuating external fields. In the present paper, we approach this goal by studying Hamiltonian and Liouvillian learning for analog quantum simulators. Previous works have studied various scenarios for Hamiltonian and Liouvillian learning [28, 29, 30, 31, 32, 33, 34, 35, 33, 34, 36, 37, 38], for instance, by comparing with a trusted simulator [28], or based on the preparation of steady states [30, 31, 32, 33, 34]. Alternative approaches are based on dynamics in (long-time) quenches [35, 33, 34, 36], or the estimation of the time-derivatives of few-qubit observables from short-time evolution [37, 38].
Here, we will be interested in Hamiltonian and Liouvillian learning from dynamics in long-time quenches of product states, with only few experimental requirements such as the preparation of product states, and product measurements. We assume that the dynamics of the experimental quantum simulator is described by a master equation, where the Hamiltonian acts as a generator of the coherent many-body dynamics, while Lindbladian terms model the noise. The goal of Hamiltonian and Liouvillian learning is to learn the operator structure, reminiscent of a principal component analysis [39], of both the many-body Hamiltonian, as one-, two- or few-body interaction terms including their couplings, and the quantum jump operators in the dissipative Liouvillian, representing local or non-local (global) quantum and classical noise. The scalability and efficiency of Hamiltonian and Liouvillian learning are related to the assumption that physical Hamiltonians and Liouvillians will only involve few-body interactions and quantum jump operators, leading to a polynomial scaling of the number of terms to be learned with system size.
@@
The paper is structured as follows. Section II outlines the specific scenario we’re examining and the core theoretical concepts involved. Section II.1 establishes some constraints on the system’s Hamiltonian and Liouvillian, measurable through simple quench experiments. We then present the main equations that enable us to infer the Hamiltonian and Liouvillian from experimental data in Sections II.2 to II.4, and the effect of shot-noise in Section II.5. Furthermore, we propose and discuss various strategies for Hamiltonian and Liouvillian learning in Section II.6. We compare different ansätze for the operator content, using an experimentally measurable quantity, which we identify as a learning error. These ansätze are derived by re-parametrizing our ansatz, typically involving data recycling and classical post-processing. The learning process can be divided into two phases as a function of experimental runs $N_{\rm runs}$: in the early phase, the learning error is dominated by shot-noise and decays $\sim N_{\rm runs}^{-1/2}$. In the later phase, systematic errors become dominant due to missing terms in the ansatz, or an insufficient ansatz. This leads to a constant learning error independent of the number of measurements, indicating the need to extend our ansatz. Finally, in Section III, we showcase learning protocols for various model scenarios through numerical simulations.
@@ II.1 Background
$\equiv$ $\mathcal{L}_{H}(\varrho)+\mathcal{L}_{D}(\varrho),$
comprising a coherent term $\mathcal{L}_{H}(\varrho)=-i[H,\varrho]$ with many-body Hamiltonian $H$ and a dissipative term $\mathcal{L}_{D}$, referred to as the Liouvillian. The Lindblad quantum jump operators $l_{k}$ describe dissipative processes coupling the system to an environment. Here, $\gamma_{k}\geq 0$ is the physical domain for the corresponding damping rates, where the dynamics is described by a completely positive and trace-preserving map [41, 42]00footnotetext: The Hamiltonian, $H$, the dissipation rates, $\gamma_{k}$, and the Lindblad operators, $l_{k}$, are uniquely determined by the dynamics, if one requires the following: (i) $H$ is traceless, (ii) the $l_{k}$ are traceless and orthonormal, i.e., $\mathrm{tr}(l_{k})=0$ and $\mathrm{tr}(l_{m}^{\dagger}l_{n})=\delta_{mn}$, and (iii) the $\gamma_{k}$ are not degenerate [41].. Throughout this work, we assume that the Hamiltonian and the Liouvillian are time-independent. In typical experimental settings, the Hamiltonian terms are significantly larger than dissipative processes $\norm{H}\gg\gamma_{k}$.
@@
The Liouvillian, $\mathcal{L}_{D}$, in Eq. (II.1) is defined by its Lindblad, or quantum jump operators $l_{k}$. Examples of Lindblad operators, that typically appear in experiments, include spontaneous emission, described by local Lindblad operators $l_{k}=\sigma^{-}_{k}$, or local dephasing, represented by $l_{k}=\sigma^{z}_{k}$. Besides local dissipation, we will be interested in identifying the presence of collective dissipative effects, for instance, in the form of collective dephasing caused by globally fluctuating laser fields or effective magnetic fields, leading to a collective Lindblad operator $l=\sum_{k}\sigma^{z}_{k}$.
@@
$\expectationvalue{H}_{T}-\expectationvalue{H}_{0}=\\ \frac{1}{2}\sum_{k}\gamma_{k}\int_{0}^{T}\expectationvalue*{l_{k}^{\dagger}[H,l_{k}]+[l_{k}^{\dagger},H]l_{k}}_{t}\mathrm{d}t.$ (4)
@@
As a final remark, let us elaborate on the experimental procedure that we are considering which can be used to probe the conditions in Eqs. (3) and (4) (see also Fig. 1 for an illustration). Starting from a product state, |$\ket{\psi(0)}=\ket{\psi_{1}}\otimes\dots\otimes\ket{\psi_{N}}$, which can be experimentally prepared with high fidelity, one evolves the state under the Hamiltonian and Liouvillian for some time $t$. The resulting state, $\varrho(t)=\exp{t(\mathcal{L}_{H}+\mathcal{L}_{D})}(\varrho(0))$ ), is measured in a product basis, for instance, in the Pauli basis. Clearly, the limiting quantity, here, and in the following, will be the total number of runs of quench experiments, which we will denote by $N_{\rm runs}$. Given, however, that many-body Hamiltonians typically consist of a few quasi-local operators, many of the required measurements can be carried out simultaneously, i.e., in a single run, using classical post-processing.
Figure 1: Our learning procedure starts by choosing an ansatz for the operator structure of the Hamiltonian and Liouvillian. The experimental procedure requires the preparation of multiple product states. These states are evolved under the unknown dynamics for different times $T\in\{T_{i}\}$ (and multiple intermediate times), followed by a measurement of a few quasi-local, few-body correlation functions, most of which can be measured simultaneously. One obtains time-discretized estimates of time-traces of observables, where times $T_{i}$ are measured using a larger measurement budget compared to intermediate times. The constraint matrix $M_{H}$ is estimated from measurements at times $T_{i}$. The constraint matrix $M_{D}$ is obtained from estimating the integrals in Eq. (8) using Simpson’s rule from measurements at intermediate times. To assess the ansatz, one considers our figure of merit $\lambda_{1}/\lambda_{2}$, or the reconstructed parameters $\bm{c}^{\rm rec}$ and $\bm{d}^{\rm rec}$, for which one can compute error bars using bootstrapping. In early stages of the learning procedure $\lambda_{1}/\lambda_{2}$ decreases with $N_{\rm runs}^{-1/2}$, indicating that the dominant source of error is shot noise. In later stages, $\lambda_{1}/\lambda_{2}$ remains constant, indicating that our ansatz is missing terms. Based on this assessment, and comparing different ansätze, we can identify an ansatz comprising the dominant operator content and few relevant parameters, resulting in an early shot-noise scaling and a low plateau. On top of that, one may also learn the overall scale of $H$, exclude conserved quantities, or learn the complete Liouvillian by including additional constraints.
II.2 Hamiltonian and Liouvillian learning
Often in, an experimental setting, the detailed structure of the Hamiltonian and the Liouvillian are unknown. We present here a method to learn the operator content of $H$ and $\mathcal{L}_{D}$, and the corresponding parameters from experimental data. For instance, in the context of the spin system in and below Eq. (II.1), identifying the operator content means identifying the Pauli operators that appear in the decomposition of $H$ and the Lindblad operators $l_{k}$.
@@ II.2 Hamiltonian and Liouvillian learning
$\mathcal{D}(\bm{d})=\frac{1}{2}\sum_{k}d_{k}\left([a_{k}\varrho,a_{k}^{\dagger}]+[a_{k},\varrho a_{k}^{\dagger}]\right),$ (6)
@@
The matrices $M_{H}$ and $M_{D}(\bm{d})$ are $p\times n$ matrices defined by $M_{H}=(M)_{ij}=\expectationvalue*{h_{j}}_{i,0}-\expectationvalue*{h_{j}}_{i,T}$, and $M_{D}(\bm{d})=\sum_{k}d_{k}M^{(k)}$, with
$[M^{(k)}]_{ij}=\int_{0}^{T}\expectationvalue*{a_{k}^{\dagger}[h_{j},a_{k}]+[a_{k}^{\dagger},h_{j}]a_{k}}_{i,t}\mathrm{d}t,$ (8)
@@
In the discussion above, we have implicitly assumed that our ansatz is chosen such as to contain the Hamiltonian and Liouvillian, i.e., there exist vectors $\bm{c}^{H}$, and $\bm{d}^{\mathcal{L}}$, such that $A(\bm{c}^{H})=H$, and $\mathcal{D}(\bm{d}^{\mathcal{L}})=\mathcal{L}_{D}$. However, in practice we do not expect this to be the case, and we will call such an ansatz insufficient  [^3] As an example, the Hamiltonian may not only contain operators $h_{j}$, but also additional operators $h_{j}^{\prime}$ not contained in our ansatz. On the level of Eq. (7) this would result in a truncation of the matrices $M_{H}$ and $M_{D}(\bm{d})$ which then contain fewer columns than would be required in order to reconstruct $\bm{c}^{H}$, or, stated differently, the remaining columns are linearly independent. Nevertheless, we can determine the parameters of our ansatz minimizing the violation of the energy conditions in Eq. (4). This amounts to minimizing the squared residuals in Eq. (7), i.e.,
$(\bm{c}^{\rm rec},\bm{d}^{\rm rec})=\operatorname*{arg\,min}_{(\bm{c},\bm{d}):\norm{\bm{c}}=1,\bm{d}\geq 0}\norm{\quantity[M_{H}+M_{D}(\bm{d})]\bm{c}},$ (9)
where $\|\bm{x}\|$ denotes the $2$-norm. Computing the minimum of the cost function in Eq. (9) corresponds to finding non-negative dissipation rates $\bm{d}^{\rm rec}$ for which the smallest singular value, $\lambda_{1}$, of the constraint matrix $M_{H}+M_{D}(\bm{d})$ attains its minimum. The corresponding $\bm{c}^{\rm rec}$ is then the right-singular vector corresponding to $\lambda_{1}$. In the following we will denote the singular vectors of the constraint matrix by $\lambda_{n}\geq\dots\geq\lambda_{2}\geq\lambda_{1}\geq 0$.
In an experiment, the constraint matrices $M_{H}$ and $M_{D}(\bm{d})$ need to be estimated from data. The matrix $M_{H}$ can be determined experimentally by measuring $\expectationvalue{h_{j}}_{0}$ and $\expectationvalue{h_{j}}_{T}$ for all input states $i$. As the $h_{j}$ are few-body operators many of them commute and can thus be measured jointly. To obtain $M_{D}$, one needs to estimate the integrals in Eq. (8). Here, we use the composite Simpson’s rule which approximates an integral as a series of parabolic segments between $N_{t}$ equally spaced points $t_{m}=m\cdot\Delta t$, with $m\in 0,\dots,N_{t}$ and $\Delta t=T/N_{t}$. It reads
$\int_{0}^{T}\expectationvalue*{h}_{t}\mathrm{d}t\approx\frac{1}{3}\Delta t\left[\expectationvalue*{h}_{0}+4\sum_{k=0}^{N_{t}/2}\expectationvalue*{h}_{t_{2k-1}}\right.\\ \left.+2\sum_{k=0}^{N_{t}/2-1}\expectationvalue*{h}_{t_{2k}}+\expectationvalue*{h}_{T}\right].$ (10)
Therefore, the estimation of $M_{D}(\bm{d})$ requires additional measurements at times $t_{m}\in[0,T]$ of operators $h_{j}$, see also Fig. 1 for an illustration. In the following, we choose a sufficiently small time step $\Delta t$ to ensure that errors arising from discretizing the integral can be disregarded in comparison to shot-noise errors  [^4] Indeed, for fixed time $T$ one obtains $\int_{0}^{T}\expectationvalue*{h}_{t}\mathrm{d}t=I(\Delta t)+K\Delta t^{4}$, where $I(\Delta t)$ is the integral approximation, and $K$ is a constant. When expressed in terms of $N_{t}$ this reads $I(N_{t})+\tilde{K}/N_{t}^{4}$. However, as $I(N_{t})$ can only be estimated from data, one obtains $I(N_{t})=\tilde{I}(N_{t})+\epsilon$, only up to a statistical error $\epsilon$. As $I(N_{t})$ is linear, the variance of $\epsilon$ roughly scales like $1/(N_{t}N_{s})$, where $N_{s}$ is the number of shots per time-point..
II.3 Learning the overall scale, and conserved quantities

### v2

@@ Abstract
We discuss Hamiltonian and Liouvillian learning for analog quantum simulation from non-equilibrium quench dynamics in the limit of weakly dissipative many-body systems. We present and compare various methods and strategies to learn the operator content of the Hamiltonian and the Lindblad operators of the Liouvillian. We compare different ansätze based on an experimentally accessible ‘learning error’ which we consider as a function of the number of runs of the experiment. Initially, the learning error decreases with the inverse square root of the number of runs, as the error in the reconstructed parameters is dominated by shot noise. Eventually the learning error remains constant, allowing us to recognize missing ansatz terms. A central aspect of our approaches is to (re-)parametrize ansätze by introducing and varying the dependencies between parameters. This allows us to identify the relevant parameters of the system, thereby reducing the complexity of the learning task. Importantly, this (re-)parametrization relies solely on classical post-processing, which is compelling given the finite amount of data available from experiments. We illustrate and compare our methods with two experimentally relevant spin models.
@@ I Introduction
An outstanding challenge in quantum simulation is the ability to predict properties of many-body observables with controlled error while scaling to a regime of potential quantum advantage [21, 22, 23, 24]. Given the increase of complexity of these systems, methods to characterize, and thus verify, the proper functioning of quantum simulators are required [25, 26, 27]. This includes verification, that the correct many-body Hamiltonians are being implemented, and a complete characterization of (weak) decoherence due to unwanted couplings to an environment or fluctuating external fields. In the present paper, we approach this goal by studying Hamiltonian and Liouvillian learning for analog quantum simulators. Previous works have studied various scenarios for Hamiltonian and Liouvillian learning [28, 29, 30, 31, 32, 33, 34, 35, 36, 34, 35, 37, 38, 39, 40], for instance, by comparing with a trusted simulator [28], or based on the preparation of steady states [31, 32, 33, 34, 35]. Alternative approaches are based on dynamics in (long-time) quenches [29, 36, 34, 35, 37], the estimation of the time-derivatives of few-qubit observables from short-time evolution [38, 39], or require additional resources such as intermediate gates [41, 42].
Here, we will be interested in Hamiltonian and Liouvillian learning from dynamics in long-time quenches, with only few experimental requirements such as the preparation of product states, and product measurements. We assume that the dynamics of the experimental quantum simulator is described by a master equation, where the Hamiltonian acts as a generator of the coherent many-body dynamics, while Lindbladian terms model the noise. The goal of Hamiltonian and Liouvillian learning is to learn the operator structure, reminiscent of a principal component analysis [43], of both the many-body Hamiltonian, as one-, two- or few-body interaction terms including their couplings, and the quantum jump operators in the dissipative Liouvillian, representing local or non-local (global) quantum and classical noise. The scalability and efficiency of Hamiltonian and Liouvillian learning are related to the assumption that physical Hamiltonians and Liouvillians will only involve few-body interactions and quantum jump operators, leading to a polynomial scaling of the number of terms to be learned with system size.
@@
The paper is structured as follows. Section II outlines the specific scenario we’re examining and the core theoretical concepts involved. Section II.1 establishes some constraints on the system’s Hamiltonian and Liouvillian, measurable through simple quench experiments. We then present the main equations that enable us to infer the Hamiltonian and Liouvillian from experimental data in Sections II.2 and II.3. The first approach is based on the equation of motion of the expectation value of general observables, known as Ehrenfest’s theorem, while the second approach is based on generalized energy conservation. In Section III, we compare the two methods and showcase learning protocols for various model scenarios through numerical simulations. We compare different ansätze for the operator content, using an experimentally measurable quantity, which we identify as a learning error. These ansätze are derived by re-parametrizing our ansatz, typically involving data recycling and classical post-processing. The learning process can be divided into two phases as a function of experimental runs $N_{\rm runs}$: in the early phase, the learning error is dominated by shot-noise and decays $\sim N_{\rm runs}^{-1/2}$. In the later phase, systematic errors become dominant due to missing terms in the ansatz, or an insufficient ansatz. This leads to a constant learning error independent of the number of measurements, indicating the need to extend our ansatz.
@@ II.1 Background
Figure 1: A quantum simulation experiment starts with a specific target Hamiltonian $H^{\rm sim}$ of interest, for instance, spin-models that we consider in this work. This Hamiltonian finds a natural realization in a quantum simulation platform that aims to simulate the dynamics of $H^{\rm sim}$ in a well controlled environment. In practice, however, the precise Hamiltonian such a device is implementing is unknown and needs to be determined. In addition, the system is typically (weakly) coupled to an environment so that the the dynamics is well described by a quantum master equation, cf. Eq. (II.1). Hamiltonian and Liouvillian learning aims to reconstruct the Hamiltonian and Liouvillian using a finite measurement budget. Here, we consider simple quench experiments with product states as inputs, and product measurements. Relevant questions that need to be answered include, but are not limited to, the existence of additional (small) terms in the Hamiltonian, as well as the operator content of the Liouvillian, i.e., the quantum jump operators and dissipation rates. This information can then be used as a feedback to the experiment, improving the quantum simulation, or to verify that indeed the correct dynamics has been implemented.
@@
$\equiv$ $\mathcal{H}(\varrho)+\mathcal{L}(\varrho),$
comprising a coherent term $\mathcal{H}(\varrho)=-i[H,\varrho]$ with many-body Hamiltonian $H$ and a dissipative term $\mathcal{L}$, referred to as the Liouvillian. The Lindblad quantum jump operators $l_{k}$ describe dissipative processes coupling the system to an environment. Here, $\gamma_{k}\geq 0$ is the physical domain for the corresponding damping rates, where the dynamics is described by a completely positive and trace-preserving map [45, 46]00footnotetext: The Hamiltonian, $H$, the dissipation rates, $\gamma_{k}$, and the Lindblad operators, $l_{k}$, are uniquely determined by the dynamics, if one requires the following: (i) $H$ is traceless, (ii) the $l_{k}$ are traceless and orthonormal, i.e., $\mathrm{tr}(l_{k})=0$ and $\mathrm{tr}(l_{m}^{\dagger}l_{n})=\delta_{mn}$, and (iii) the $\gamma_{k}$ are not degenerate [45].. Throughout this work, we assume that the Hamiltonian and the Liouvillian are time-independent. In typical experimental settings, the Hamiltonian terms are significantly larger than dissipative processes $\norm{H}\gg\gamma_{k}$.
@@
The Liouvillian, $\mathcal{L}$, in Eq. (II.1) is defined by its Lindblad, or quantum jump operators $l_{k}$. Examples of Lindblad operators that typically appear in experiments include spontaneous emission, described by local Lindblad operators $l_{k}=\sigma^{-}_{k}$, or local dephasing, represented by $l_{k}=\sigma^{z}_{k}$. Besides local dissipation, we will be interested in identifying the presence of collective dissipative effects, for instance, in the form of collective dephasing caused by globally fluctuating laser fields or effective magnetic fields, leading to a collective Lindblad operator $l=\sum_{k}\sigma^{z}_{k}$.
@@
$\expectationvalue{O}_{T}-\expectationvalue{O}_{0}=-i\int_{0}^{T}\expectationvalue{\quantity[O,H]}\mathrm{d}t\\ +\frac{1}{2}\sum_{k}\gamma_{k}\int_{0}^{T}\expectationvalue*{l_{k}^{\dagger}[O,l_{k}]+[l_{k}^{\dagger},O]l_{k}}_{t}\mathrm{d}t,$ (4a)
@@
$\expectationvalue{H}_{T}-\expectationvalue{H}_{0}=\\ \frac{1}{2}\sum_{k}\gamma_{k}\int_{0}^{T}\expectationvalue*{l_{k}^{\dagger}[H,l_{k}]+[l_{k}^{\dagger},H]l_{k}}_{t}\mathrm{d}t.$ (4b)
It describes generalized energy conservation, including the loss of total energy of the system during a quench of duration $T$. If $\gamma_{k}=0$ for all $k$, this equation indicates the conservation of energy. In our Hamiltonian and Liouvillian learning protocol, Eqs. (4a) and (4b) will play a crucial role. We will generalize the protocol of Ref. [36] for Hamiltonian learning in the absence of dissipation, and present it in a form particularly suited for learning $H$ and $\cal L$ in the limit of weak dissipation.
[…]
II.2 Hamiltonian and Liouvillian learning from the generalized Ehrenfest theorem
Often, in an experimental setting, the detailed structure of the Hamiltonian and the Liouvillian are unknown. We present here a method to learn the operator content of $H$ and $\mathcal{L}$, and the corresponding parameters from experimental data. For instance, in the context of the spin system in and below Eq. (II.1), identifying the operator content means identifying the Pauli operators that appear in the decomposition of $H$ and the Lindblad operators $l_{k}$.
@@ II.2 Hamiltonian and Liouvillian learning from the generalized Ehrenfest theorem
$\mathcal{D}(\bm{d})=\frac{1}{2}\sum_{k=1}^{m}d_{k}\left([a_{k}\varrho,a_{k}^{\dagger}]+[a_{k},\varrho a_{k}^{\dagger}]\right),$ (6)
with Lindblad operators $\{a_{k}\}$, and parameterized by the corresponding non-negative dissipation rates, $\bm{d}=(d_{1},d_{2},\ldots,d_{m})$. Here, $a_{k}$ is intended to be an ansatz operator for a single Lindblad operator, $l_{k}$. However, this must not be the case as we will discuss in Section II.3.
@@
$K_{H}\bm{c}+K_{D}\bm{d}=\bm{b}.$ (7)
Here, $K_{H}$ is a $p\times n$ matrices, and $K_{D}$ is a $p\times m$ matrix, with entries defined by
$(K_{H})_{ij}=-i\int_{0}^{T}\expectationvalue*{[O_{i},h_{j}]}_{t}\mathrm{d}t,$ (8)
and
$(K_{D})_{ik}=\frac{1}{2}\int_{0}^{T}\expectationvalue*{a_{k}^{\dagger}[O_{i},a_{k}]+\mathrm{h.c.}}_{t}\,\mathrm{d}t,$ (9)
and the vector $\bm{b}$ is defined by $b_{i}=\expectationvalue*{O_{i}}_{T}-\expectationvalue*{O_{i}}_{0}$. As these equations hold for any density matrix, $\varrho$, our protocol, similar to the Hamiltonian learning protocol in Ref. [36], is resistant to state-preparation errors. Moreover, although we have imposed the above equations for multiple observables, $\{O_{i}\}$, and a single input state, $\varrho$, one may also consider multiple input states. Note also, that similar constraints have also been used in Ref. [33] to learn Liouvillians from their steady-states.
In the discussion above, we have implicitly assumed that our ansatz is chosen such as to contain the Hamiltonian and Liouvillian, i.e., there exist vectors $\bm{c}^{H}$, and $\bm{d}^{\gamma}$, such that $A(\bm{c}^{H})=H$, and $\mathcal{D}(\bm{d}^{\gamma})=\mathcal{L}$. However, in practice we do not expect this to be the case, and we will call such an ansatz insufficient  [^2] As an example, the Hamiltonian may not only contain operators $h_{j}$, but also additional operators $h_{j}^{\prime}$ not contained in our ansatz. On the level of Eq. (11) this would result in a truncation of the matrix $M$ which then contains fewer columns than would be required in order to reconstruct $\bm{c}^{H}$. Nevertheless, we can determine the parameters of our ansatz minimizing the violation of Eq. (7) by minimizing the squared residuals;
$(\bm{c}^{\rm rec},\bm{d}^{\rm rec})=\operatorname*{arg\,min}_{(\bm{c},\bm{d}):\bm{d}\geq 0}\norm{K_{H}\bm{c}+K_{D}\bm{d}-\bm{b}},$ (10)
where $\|\bm{x}\|$ denotes the $2$-norm. In case of an insufficient ansatz, the value of the linear optimization problem, $\min_{(\bm{c},\bm{d})}\norm{K_{H}\bm{c}+K_{D}\bm{d}-\bm{b}}$, will be strictly bounded away from zero.
Figure 2: Hamiltonian and Liouvillian learning begins by choosing an ansatz for the operator structure of the Hamiltonian and Liouvillian. Then, one measures time-resolved expectation values of correlation functions, and estimates of their integrals over a period of total time $T$. Here, we choose to measure the correlation functions at discrete points and estimate the integral using Simpson’s rule. The learning progress is assessed by considering the ratio $\lambda_{1}/\lambda_{2}$, which can be directly obtained from experimental data, as a function of $N_{\rm runs}$. This allows us to compare different ansätze, and to notice when terms in the ansatz are missing. changing the ansatz involves reparametrization, which typically can be done by classical post-processing.
II.3 Hamiltonian and Liouvillian learning from generalized energy conservation
Instead of inserting our ansatz into Eq. (4a), we may also insert the ansatz into the much simpler generalized energy conservation condition in Eq. (4b). Imposing the resulting constraint for a set of initial (product) states $\{\varrho_{i}\}$, where $i=1,\dots,p$, leads to the following simple matrix equation;
@@ II.3 Hamiltonian and Liouvillian learning from generalized energy conservation
The matrices $M_{H}$ and $M_{D}(\bm{d})$ are $p\times n$ matrices defined by
$(M_{H})_{ij}=\expectationvalue*{h_{j}}_{i,0}-\expectationvalue*{h_{j}}_{i,T},$ (12)
and $M_{D}(\bm{d})=\frac{1}{2}\sum_{k}d_{k}M^{(k)}$, with
$(M^{(k)})_{ij}=\int_{0}^{T}\expectationvalue*{a_{k}^{\dagger}[h_{j},a_{k}]+[a_{k}^{\dagger},h_{j}]a_{k}}_{i,t}\mathrm{d}t,$ (13)
@@
Similar as above, in the case of an insufficient ansatz that cannot fulfill Eq. (11), one can reconstruct those parameters minimizing the violation of generalized energy conservation;
$(\bm{c}^{\rm rec},\bm{d}^{\rm rec})=\operatorname*{arg\,min}_{(\bm{c},\bm{d}):\norm{\bm{c}}=1,\bm{d}\geq 0}\norm{\quantity[M_{H}+M_{D}(\bm{d})]\bm{c}}.$ (14)
Without dissipation, where $\bm{d}=0$, this is a linear optimization problem, while in the presence of dissipation, i.e., $\bm{d}>0$, Eq. (14) involves non-linear optimization over $\bm{c}$ and $\bm{d}$. Note, that for a fixed $\bm{d}$, the minimization over $\bm{c}$ results in the smallest singular value, $\lambda_{1}$, of the constraint matrix $M_{H}+M_{D}(\bm{d})$. Computing the minimum of the cost function in Eq. (14) then corresponds to finding non-negative dissipation rates $\bm{d}^{\rm rec}$ for which the smallest singular value, $\lambda_{1}$, of the constraint matrix $M_{H}+M_{D}(\bm{d})$ attains its minimum. Numerically we find $\bm{d}^{\rm rec}$ using the "Dividing Rectangles" (DIRECT) algorithm implemented in SciPy with default settings  [^3] The lower bound for each $d_{k}$ is set to $d_{k}^{\mathrm{min}}=0$, and the upper bound is set proportional to the maximum dissipation rate of the model $d_{k}^{\mathrm{max}}=k\gamma_{\mathrm{max}}=k\mathrm{max}(\{\gamma_{k}\})$. We check convergence for different values of $k$ and different numbers of cost-function evaluations.. The corresponding $\bm{c}^{\rm rec}$ is then the right-singular vector corresponding to $\lambda_{1}$. In the following we will denote the singular values of the constraint matrix by $\lambda_{n}\geq\dots\geq\lambda_{2}\geq\lambda_{1}\geq 0$.
II.3.1 Learning the overall scale, and conserved quantities